2) Calculate the percent ionization of 0.130 M lactic acid in a solution containing 8.5×10−3  M sodium lactate.

Express the percent ionization to two significant digits.
Answer #1

    HC3H5O3(aq) --------------> C3H5O3^- (aq) + H^+ (aq)

PKa = 3.86

PH   = Pka + log[NaC3H5O3]/[HC3H5O3]

PH   = 3.86 + log(8.5*10^-3/0.13)

PH   = 3.86 -1.1845

PH = 2.6755

-log[H^+] = 2.6755

[H^+]   = 10^-2.6755   = 0.0021M

percent ionization = [H^+]*100/concentration of acid

                           = 0.0021*100/0.13

                           = 1.62%